An exclusive Yoghurt manufacturer sells 4,000 gallons per month at a price of GHS 40

each. When the price is reduced to GHS 30 sales increase to 6,000 gallons per month.

a. Calculate the price elasticity of demand for the Yoghurts over this price range.

b. Is demand elastic, unit elastic or inelastic?

c. Calculate the change in revenue due to the change in price.


1
Expert's answer
2021-05-03T10:55:23-0400

Solution:

a.). Price elasticity of demand (PED) ="\frac{\%\;change\; in\; quantity\; demanded} {\%\; change\; in\; price}"


% change in qty demanded = "\frac{Q_{2} -Q_{1}}{(Q_{2}+Q_{1})/2 } \times 100 = \frac{6000 -4000}{(6000+4000)/2 } \times 100"


= "\frac{2000}{5000 } \times 100 = 40\%"


% change in price = "\frac{P_{2} -P_{1}}{(P_{2}+P_{1})/2 } \times 100 = \frac{30 -40}{(30+40)/2 } \times 100"


"= \frac{-10}{33 } \times 100 = -28.57\%"


Price elasticity of demand (PED) = "\frac{40\%}{-28.57\% } = -1.40"

PED = 1.40

 

b.). The demand is elastic. This is because PED is greater than one, meaning that consumers are very sensitive to Yoghurt price changes.

 

c.). Revenue = Price per unit "\times" the number of units sold (P "\times" Q)

Initial revenue before the price change = 40 "\times" 4000 = GHS160,000

Revenue after the price change = 30 "\times" 6000 = GHS180,000

Change in revenue due to change in price = 180,000 – 160,000 = 20,000

Change in revenue due to change in price = GHS20,000